OVHCloud Bare Metal Cloud Status

Current status
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#6932 — vps-20-n5 et vps-21-n5
Scheduled Maintenance Report for Bare Metal Cloud
We are going to update the switches managing the offers VPS and PCI.
The upgrade will be done on 2 ways independently one after the other. There is no expected interruption.

We planned the update in the night of Wednesday 27th June to Thursday 28th June at 00:00 (CEST).


Date: 2012-07-26 03:16:48 UTC
Both switches have been updated and are running.

Date: 2012-07-26 01:19:15 UTC
The traffic has been switched.

We are updating vps-21-n5.

Date: 2012-07-26 01:18:38 UTC
The switch is updated.

We are switching the traffic on the updated switch vps-20-n5.

Date: 2012-07-26 00:23:45 UTC
Traffic is now switched.

We will update vps-20-n5.

Date: 2012-07-26 00:23:12 UTC
We are switching traffic on vps-21-n5.

Date: 2012-07-24 14:37:33 UTC
We detect a bug in the version of the 2 switchs.

2012 Jul 24 07:10:40 vps-20-n5 %ETHPORT-5-IF_SEQ_ERROR: Error (\"sequence timeout\") communicating with MTS_SAP_STP for opcode MTS_OPC_ETHPM_PORT_LOGICAL_CLEANUP (RID_PORT: Ethernet107/1/18) - Ethpm didn't receive a Link UP from the driver on time or had an
internal error, check Port Client and Ethpm logs and debug

We will proceed to their update during the night of the 25th of July till the 26th at 00:00.
The infrastructure as it is redundant, no interruption of traffic is expected.

Date: 2012-06-28 11:33:55 UTC
The switch vps-20-n5 is updated.

Everything is operational.

Date: 2012-06-28 11:33:02 UTC
All the trafic is switched on vps-21-n5.

We are updating the switch vps-20-n5 with its Fex.

Date: 2012-06-28 11:30:56 UTC
The update was done with success. All the trafic is on vps-20-n5.

The switch vps-21-n5 is now updated and operational.

We switch the vps-20-n5 on vps-21-n5.

Date: 2012-06-28 11:29:06 UTC
We start with the switch vps-21-n5.
Posted Jun 28, 2012 - 11:27 UTC